### v2

@@ 7 Numerical illustration of PICSE
To describe the competitors of PICSE, suppose $S=1/n\sum_{i=1}^{n}y_{i}y_{i}^{\top}$ is a sample covariance matrix of the data with $y_{i}=\text{vec}(Y_{i})$. Let $\tilde{K}$ and $\tilde{C}$ be the separable and the core components of $S$, respectively. Then we consider the Kronecker MLE (KMLE) [soloveychik2016, srivastava2008, dutilleul999], which is exactly $\tilde{K}$, and the core shrinkage estimator (CSE) proposed by [hoff2023]. The CSE is defined by $\tilde{K}^{1/2}((1-\hat{w})\tilde{C}+\hat{w}I_{p})\tilde{K}^{1/2}$, where $\tilde{K}^{1/2}$ is a symmetric square root of $\tilde{K}$ and the shrinkage amount $\hat{w}$ is estimated via empirical Bayes (see Section $3.1$ of [hoff2023]). Additionally, we also consider the baseline methods for PICSE based on the initialization in (35). Namely, we obtain the initial estimate of the population covariance by plugging the initialization in (35) into (29), denoted Base-AI ($K^{1/2}\in\mathcal{S}_{p_{1},p_{2}}^{++}$) and Base-Chol $(K^{1/2}\in\mathcal{L}_{p_{1},p_{2}}^{++})$. Accordingly, we consider two versions of PICSE, denoted by PI-AI ($K^{1/2}\in\mathcal{S}_{p_{1},p_{2}}^{++}$) and PI-Chol $(K^{1/2}\in\mathcal{L}_{p_{1},p_{2}}^{++})$. The baseline methods are considered to examine whether leveraging the geometry of $\mathcal{C}_{p_{1},p_{2},r}$ to find the optimal direction in updating $A$ leads to the improved estimate.
@@
$(p_{1},p_{2},r)$ Method $n=p/8$ $n=p/4$ $n=p/2$ $n=p$ $n=2p$
$(16,10,4)$ KMLE 0.964 * * * *
CSE 0.607 0.527 0.460 0.365 0.271
Base-AI 0.596 0.392 0.279 0.192 0.132
Base-Chol 0.627 0.412 0.293 0.201 0.138
PI-AI 0.541 0.312 0.206 0.139 0.095
PI-Chol 0.555 0.324 0.217 0.146 0.101
$(12,8,4)$ KMLE 0.939 * * * *
CSE 0.703 0.608 0.511 0.402 0.301
Base-AI 0.839 0.542 0.373 0.256 0.177
Base-Chol 0.884 0.569 0.389 0.267 0.185
PI-AI 0.940 0.486 0.301 0.203 0.136
PI-Chol 0.860 0.500 0.312 0.211 0.144
[…]
$(p_{1},p_{2},r)$ Method $n=p/8$ $n=p/4$ $n=p/2$ $n=p$ $n=2p$
$(16,10,4)$ KMLE 0.910 0.910 0.910 0.910 0.910
CSE 0.668 0.534 0.477 0.377 0.279
Base-AI 0.762 0.490 0.354 0.245 0.169
PI-AI 0.598 0.354 0.248 0.173 0.120
PI-Chol 0.587 0.352 0.249 0.174 0.123
$(12,8,4)$ KMLE 0.872 0.871 0.871 0.869 0.869
CSE 0.803 0.580 0.484 0.385 0.287
Base-AI 0.917 0.601 0.422 0.303 0.206
PI-AI 0.867 0.482 0.322 0.226 0.156
PI-Chol 0.852 0.469 0.323 0.227 0.158
[…]
8 Concluding remark
@@
We have studied the differential geometry of the fixed-rank core covariance manifold $\mathcal{C}_{p_{1},p_{2},r}^{+}$ when $p_{1}/p_{2}+p_{2}/p_{1}<r\leq p$. When $r<p$, we established that $\mathcal{C}_{p_{1},p_{2},r}^{+}$ is a smooth manifold after removing the set of canonically decomposable matrices. For the full-rank case, we further established a diffeomorphic relationship between $\mathcal{S}_{p}^{++}$ and $\mathcal{S}_{p_{1},p_{2}}^{++}\times\mathcal{C}_{p_{1},p_{2}}^{++}$, providing a new insight into the smooth structure of $\mathcal{S}_{p}^{++}$ in terms of the separability. We also derived differential quantities on $\mathcal{C}_{p_{1},p_{2},r}^{+}$, including tangent vectors, the differentials of the diffeomorphism when $r=p$, and Riemannian gradient and Hessian operator on $\mathcal{C}_{p_{1},p_{2},r}$, and $\mathcal{C}_{p_{1},p_{2}}^{++}$ under the Euclidean metric, with respect to which $\mathcal{C}_{p_{1},p_{2}}^{++}$ is flat. The corresponding Riemannian gradient and Hessian operator are also obtained for $\mathcal{C}_{p_{1},p_{2},r}/\mathcal{O}_{r}$ via quotient geometry.
@@
An interesting future direction is to identify a Riemannian metric on $\mathcal{C}_{p_{1},p_{2},r}^{+}$ that induces nice geometric properties, such as completeness, closed-form geodesics, or nonpositive sectional curvature. One approach is to construct a group-invariant metric (see [thanwerdas2023b]). By the linear system that defines the core as in (3), the Kronecker orthogonal group $\mathcal{O}_{p_{1},p_{2}}$ acts smoothly on $\mathcal{C}_{p_{1},p_{2},r}^{+}$ via the action $(O_{2}\otimes O_{1},C)\in\mathcal{O}_{p_{1},p_{2}}\times\mathcal{C}_{p_{1},p_{2},r}^{+}\rightarrow(O_{2}\otimes O_{1})C(O_{2}\otimes O_{1})^{\top}\in\mathcal{C}_{p_{1},p_{2},r}^{+}$. However, this action is not transitive, and the invariant metric is hence not unique under this action. Thus, the metric will vary across orbits, leading to infinitely many invariant metrics. We leave this as a future direction to further explore the geometry of $\mathcal{C}_{p_{1},p_{2},r}^{+}$.
@@
We introduced the partial isotropy core shrinkage estimator (PICSE), assuming that the population core has a partial-isotropy structure. Since the partial-isotropy (factor) covariance model is often used for vector data and the covariance of a random matrix is defined by that of its vectorization (see (1)), one might ask whether PICSE can be used to estimate a factor-type covariance matrix for general $p$-dimensional vector data with correctly specified $p_{1}$ and $p_{2}$. Although technically possible, we do not recommend using PICSE for vector data as it will lose interpretation. As discussed in Section 6, the partial-isotropy core covariance model aims to make an effective departure from the separability assumption commonly used in modeling matrix-variate data. Thus, using PICSE is meaningful only when the separability assumption is reasonable. Because the assumption enables a separate inference of the correlation structures of the row and column variables [dawid1981, wang2022], the data must have two different modes for the assumption, which is not the case for general vector data.
Appendix
In Appendix, we provide the omitted proofs from the main text (Appendix A), the formulas of Euclidean derivative and Hessian operator of the negative log-likelihood $\ell$ defined in (28) to compute PICSE (Appendix B), and the additional figures that support the results in Section 7 (Appendix C).
@@ Appendix A Deferred Proofs
All the omitted proofs from the main text are provided in this section.
@@ A.1 Proofs of the Results from Section 2
Proof A.1 (Proof of Example 2.3).
The proof is based on Proposition $3$ of [hoff2023]. Suppose $K=\Sigma_{2}\otimes\Sigma_{1}$ is a Kronecker MLE of $FF^{\top}$. Then since $K^{-1/2}FF^{\top}K^{-1/2,\top}$ is a core, Proposition $3$ of [hoff2023] implies that
(36) $\sum_{(i,j)}E_{ij}\Omega_{2}E_{ij}^{\top}$ $=2\Sigma_{1},\;\sum_{(i,j)}E_{ij}^{\top}\Omega_{1}E_{ij}=2\Sigma_{2},$
@@
Since $\Sigma_{2}\in\mathcal{S}_{2}^{++}$, we have that $\sigma_{2,22}/(\sigma_{2,11}+\sigma_{2,22})=1$, which is not true unless $\sigma_{2,11}=0$. Hence, this contradicts the existence of the Kronecker MLE $K$.
Proof A.2 (Proof of Proposition 2).
@@
(37) $\begin{split}\gamma^{\prime}(t)&=\lfloor V\rfloor+\mathbb{D}(V)\exp\left(t\mathbb{D}(V)\mathbb{D}(L)^{-1}\right),\\ \gamma^{\prime\prime}(t)&=\mathbb{D}(V)^{2}\mathbb{D}(L)^{-1}\exp\left(t\mathbb{D}(V)\mathbb{D}(L)^{-1}\right).\end{split}$
@@
(38) $\begin{split}v^{\prime}(t)&=\textsf{tr}\left(\nabla f(\gamma(t))^{\top}\gamma^{\prime}(t)\right),\\ v^{\prime\prime}(t)&=\textsf{tr}\left(\nabla^{2}f(\gamma(t))[\gamma^{\prime}(t)]^{\top}\gamma^{\prime}(t)\right)+\textsf{tr}\left(\nabla f(\gamma(t))^{\top}\gamma^{\prime\prime}(t)\right).\end{split}$
@@ A.2 Proofs of the Results from Section 3
Proof A.4 (Proof of Proposition 4).
Following the main idea outlined in Section 3, it suffices to verify (14). Note that $a^{\top}a=p$, $A_{R}=p_{2}I_{p_{1}}$, and $A_{C}=p_{1}I_{p_{2}}$. By Theorem $3.1$ of [magnus1979], $K_{(q,q)}$ is a symmetric matrix whose eigenvalues are either $1$ or 1-1, with respective multiplicities $q(q+1)/2$ and q(q1)/2q(q-1)/2. Moreover, the eigenspace of $K_{(q,q)}$ corresponding to the eigenvalue $1$ (resp. 1-1) is exactly the vectorization of $\mathcal{S}_{q}$ (resp. $\textsf{Skew}_{q}$). By vec-Kronecker identity,
@@
(39) $(V-\textsf{tr}\left(V\right)/p_{1}I_{p_{1}})A_{i}=A_{i}(W-\textsf{tr}\left(W\right)/p_{2}I_{p_{2}}).$
@@
(40) $\tilde{\Lambda}_{V}\tilde{A}_{i}=\tilde{A}_{i}\tilde{\Lambda}_{W}$
where $\tilde{\Lambda}_{V}=\Lambda_{V}-\textsf{tr}\left(\Lambda_{V}\right)/p_{1}I_{p_{1}}$ and $\tilde{\Lambda}_{W}=\Lambda_{W}-\textsf{tr}\left(\Lambda_{W}\right)/p_{2}I_{p_{2}}$. The equation (40) holds for each $i$ if and only if $((\tilde{\Lambda}_{V})_{jj}-(\tilde{\Lambda}_{W})_{kk})(\tilde{A}_{i})_{jk}=0$ for any $(j,k)\in[p_{1}]\times[p_{2}]$ and $i\in[p]$. Note that $\tilde{A}_{1},\ldots,\tilde{A}_{p}$ are linearly independent by the definition of $\mathcal{H}_{p_{1},p_{2}}$. Thus, for any fixed $(j,k)\in[p_{1}]\times[p_{2}]$, there must exist at least one $i\in[p]$ for which $(\tilde{A}_{i})_{jk}$ is nonzero, otherwise the linear independence is violated. Hence, $(\tilde{\Lambda}_{V})_{jj}=(\tilde{\Lambda}_{W})_{kk}$ for any $j,k$. As such, the diagonal entries of $\tilde{\Lambda}_{V}$ and $\tilde{\Lambda}_{W}$ are all equal to some constant $c$. However, as the traces of these matrices are the same as 0, both $\tilde{\Lambda}_{V}$ and $\tilde{\Lambda}_{W}$ are zero diagonal matrices so that $V=\Lambda_{V}=c_{1}I_{p_{1}}$ and $W=\Lambda_{W}=c_{2}I_{p_{2}}$ for some constants $c_{1},c_{2}$, leading to zero vectors $J_{1}^{\top}v$ and $J_{2}^{\top}w$. Hence, $C(J_{1}^{\top})\cap C(J_{2}^{\top})=\left\{0\right\}$, concluding the proof.
@@
as $r>\alpha/\beta+\beta/\alpha\geq 2$ and $\alpha,\beta>2$, we conclude that the maximum of $m$ is strictly smaller than $r\alpha\beta$.
Proof A.6 (Proof of Proposition 7).
@@
(41) $\mathcal{P}_{U_{1}}A_{i}(I-\mathcal{P}_{W_{1}})=0,\quad(I-\mathcal{P}_{U_{1}})A_{i}\mathcal{P}_{W_{1}}=0.$
Recall that the orthogonal projection into a $d$-dimensional linear subspace of n corresponds to a unique element in a projective variety $\text{Gr}(d,n)$ as every linear subspace has its unique orthogonal projection. Namely, suppose $Z$ denotes a $d\times n$ matrix whose rows denote the basis of a $d$-dimensional subspace. Then the orthogonal projection onto this subspace is given by $Z^{\top}\textsf{adj}(ZZ^{\top})Z/|ZZ^{\top}|$, where $\textsf{adj}(M)$ is an adjoint of a square matrix $M$. By Theorem $2.1$ of [devriendt2024], each entry of $Z^{\top}\textsf{adj}(ZZ^{\top})Z$ and $|ZZ^{\top}|$ can be expressed as a quadratic polynomial in Plüker coordinates. Note that $|ZZ^{\top}|\neq 0$. Hence, multiplying the quadratic polynomial corresponding to $|ZZ^{\top}|$ in Plüker coordinates to both hand sides of the equations in (41), we see that (41) induces the system of finitely many equations of polynomials in Plüker coordinates and affine coordinates (the entries of $A_{1},\ldots,A_{r}$). Hence, if $\tilde{\mathcal{V}}_{p_{1},p_{2},r}^{(a,b)}$ is a subset of $\mathcal{Z}:=\real\mathbb{P}^{\binom{p_{1}}{a}-1}\times\real\mathbb{P}^{\binom{p_{2}}{b}-1}\times\real^{pr}$ for which (41) is satisfied, then $\tilde{\mathcal{V}}_{p_{1},p_{2},r}^{(a,b)}$ is Zariski-closed as it is exactly the zero set of finitely many polynomials over $\mathcal{Z}$. Furthermore, since there are $ab+(p_{1}-a)(p_{2}-b)$ free coordinates for each $A_{i}$ with fixed $\mathcal{P}_{U_{1}},\mathcal{P}_{W_{1}}$,

### v3

@@ 7 Illustration of PICSE
To describe the competitors of PICSE, suppose $S$ is a sample covariance matrix based on random matrices $Y_{1},\ldots,Y_{n}$. Let $\tilde{K}$ and $\tilde{C}$ be the separable and the core components of $S$, respectively. Then we consider the Kronecker MLE (KMLE) [61, 62, 25], which is exactly $\tilde{K}$, and the core shrinkage estimator (CSE) proposed by [33]. The CSE is defined by $\tilde{K}^{1/2}((1-\hat{w})\tilde{C}+\hat{w}I_{p})\tilde{K}^{1/2}$, where $\tilde{K}^{1/2}$ is a symmetric square root of $\tilde{K}$ and the shrinkage amount $\hat{w}$ is estimated via empirical Bayes (see Section $3.1$ of [33]). Additionally, we consider the baseline methods for PICSE based on the initialization in (35). Namely, we obtain the initial estimate of the population covariance by plugging the initialization in (35) into (29), denoted Base-AI ($K^{1/2}\in\mathcal{S}_{p_{1},p_{2}}^{++}$) and Base-Chol $(K^{1/2}\in\mathcal{L}_{p_{1},p_{2}}^{++})$. Accordingly, we consider two versions of PICSE, denoted by PI-AI ($K^{1/2}\in\mathcal{S}_{p_{1},p_{2}}^{++}$) and PI-Chol $(K^{1/2}\in\mathcal{L}_{p_{1},p_{2}}^{++})$. The baseline methods are considered to examine whether leveraging the geometry of $\mathcal{C}_{p_{1},p_{2},r}$ to find the optimal direction in updating $A$ leads to the improved estimate.
[…]
Figures 14 show the box plots of the relative differences across $100$ iterations with different $(p_{1},p_{2},r)$, $\lambda$, and $n$. In general, one can verify that KMLE performs poorly compared to other methods and shows a small standard deviation of the relative norms. This is because the dimension of $\mathcal{S}_{p_{1},p_{2}}^{++}$ is much lower than that of the space where the partial-isotropy core covariance lies. Thus, KMLE is already close to the pseudo-true parameter, namely, the separable component of $\Sigma$, $K$. Hence, KMLE may be estimating $K$ well but yield a poor estimate of $\Sigma$ as its core component is fixed as $I_{p}$. This implies that the Kronecker MLE is not a good estimate if the true covariance is not separable.
On the other hand, the other methods tend to show the improved performance as $n$ grows for each choice of $\lambda$ and $(p_{1},p_{2},r)$ under the models (M1)(M2). Note that both PI-AI and PI-Chol perform better than CSE and Base-AI, particularly for the small values of $n$ and $\lambda$ under both (M1)(M2). This illustrates the robustness of PICSE, allowing a slight departure from the true covariance model in (M1). Also, the performance gap between Base-AI and PICSE is more obvious with small $n$ and $\lambda$. This is because PICSE leverages the curvature of the negative log-likelihood in (28) to find the optimal $A$. Moreover, as can be seen from Figures 512 in Appendix C, while there might be no significant improvement for PICSE in estimating $K$ compared to other methods, there is in estimating $C$, particularly with small $n$ and $\lambda$. This implies that PICSE is effective in a high-dimensional regime when the true core exhibits a low-dimensional feature and is far from the separability.
Lastly, note that the only scenario where CSE performs better than PICSE and Base-AI is for small $n$ but large $\lambda$. That is, when the sample size is small and the true covariance is close to separability, CSE can perform better than these two estimators. The reason is that CSE estimates the non-spiked eigenvalue $\lambda$ via empirical Bayes and tends to shrink more toward the separability compared to PICSE. Hence, it may be less prone to overfitting for small sample sizes when $\Sigma$ is close to separability. This is further supported by the estimates of $\lambda$ for CSE and PICSE in Tables 23 under the model (M1) from Appendix C.
@@
8 Concluding remarks
@@
We have studied the geometry of the fixed-rank core covariance manifold $\mathcal{C}_{p_{1},p_{2},r}^{+}$ with $p_{1}/p_{2}+p_{2}/p_{1}<r\leq p$. When $r<p$, we established that $\mathcal{C}_{p_{1},p_{2},r}^{+}$ is a smooth manifold after removing the set of canonically decomposable matrices. For the full-rank case, we further established a diffeomorphic relationship between $\mathcal{S}_{p}^{++}$ and $\mathcal{S}_{p_{1},p_{2}}^{++}\times\mathcal{C}_{p_{1},p_{2}}^{++}$, providing a new insight into the smooth structure of $\mathcal{S}_{p}^{++}$ in terms of the separability. We also derived differential quantities on $\mathcal{C}_{p_{1},p_{2},r}^{+}$, including tangent vectors, the differentials of the diffeomorphism when $r=p$, and Riemannian gradient and Hessian operator on $\mathcal{C}_{p_{1},p_{2},r}$, and $\mathcal{C}_{p_{1},p_{2}}^{++}$ under the Euclidean metric, with respect to which $\mathcal{C}_{p_{1},p_{2}}^{++}$ is flat. The corresponding Riemannian gradient and Hessian operator are also obtained for $\mathcal{C}_{p_{1},p_{2},r}/\mathcal{O}_{r}$ via quotient geometry.
@@
An interesting future direction is to identify a Riemannian metric on $\mathcal{C}_{p_{1},p_{2},r}^{+}$ that induces nice geometric properties, such as completeness, closed-form geodesics, or nonpositive sectional curvature. One approach is to construct a group-invariant metric (see [68]). By the linear system that defines the core as in (4), the Kronecker orthogonal group $\mathcal{O}_{p_{1},p_{2}}$ acts smoothly on $\mathcal{C}_{p_{1},p_{2},r}^{+}$ via the action $(O_{2}\otimes O_{1},C)\in\mathcal{O}_{p_{1},p_{2}}\times\mathcal{C}_{p_{1},p_{2},r}^{+}\rightarrow(O_{2}\otimes O_{1})C(O_{2}\otimes O_{1})^{\top}\in\mathcal{C}_{p_{1},p_{2},r}^{+}$. However, this action is not transitive, and the invariant metric is hence not unique under this action. Thus, the metric will vary across orbits, leading to infinitely many invariant metrics. We leave this as a future direction to further explore the geometry of $\mathcal{C}_{p_{1},p_{2},r}^{+}$.
@@
We introduced the partial isotropy core shrinkage estimator (PICSE), assuming that the population core has a partial-isotropy structure. Since the partial-isotropy (factor) covariance model is often used for vector data and the covariance of a random matrix is defined by that of its vectorization (see (1)), one might ask whether PICSE can be used to estimate a factor-type covariance matrix for general $p$-dimensional vector data with correctly specified $p_{1}$ and $p_{2}$. Although technically possible, we do not recommend using PICSE for vector data as it will lose interpretation. As discussed in Section 6, the partial-isotropy core covariance model aims to make an effective departure from the separability assumption commonly used in modeling matrix-variate data. Thus, using PICSE is meaningful only when the separability assumption is valid. Because the assumption enables a separate inference of the correlation structures of the row and column variables [18, 69], the data must have two different modes for the assumption, which is not the case for general vector data.
@@ Appendix A Deferred proofs
In this section, we provide the omitted proofs from the main text.
@@ A.1 Proofs of the results from Section 2
Proof of Example 2.1.
The proof is based on Proposition $3$ of [33]. Suppose $K=\Sigma_{2}\otimes\Sigma_{1}$ is a Kronecker MLE of $FF^{\top}$. Then since $K^{-1/2}FF^{\top}K^{-1/2,\top}$ is a core, Proposition $3$ of [33] implies that
$\sum_{(i,j)}E_{ij}\Omega_{2}E_{ij}^{\top}$ $=2\Sigma_{1},\;\sum_{(i,j)}E_{ij}^{\top}\Omega_{1}E_{ij}=2\Sigma_{2},$ (36)
@@
Since $\Sigma_{2}\in\mathcal{S}_{2}^{++}$, we have that $\sigma_{2,22}/(\sigma_{2,11}+\sigma_{2,22})=1$, which is not true unless $\sigma_{2,11}=0$. Hence, this contradicts the existence of the Kronecker MLE $K$. ∎
Proof of Proposition 2.3.
@@
$\begin{split}\gamma^{\prime}(t)&=\lfloor V\rfloor+\mathbb{D}(V)\exp\left(t\mathbb{D}(V)\mathbb{D}(L)^{-1}\right),\\ \gamma^{\prime\prime}(t)&=\mathbb{D}(V)^{2}\mathbb{D}(L)^{-1}\exp\left(t\mathbb{D}(V)\mathbb{D}(L)^{-1}\right).\end{split}$ (37)
@@
$\begin{split}v^{\prime}(t)&=\textsf{tr}\left(\nabla f(\gamma(t))^{\top}\gamma^{\prime}(t)\right),\\ v^{\prime\prime}(t)&=\textsf{tr}\left(\nabla^{2}f(\gamma(t))[\gamma^{\prime}(t)]^{\top}\gamma^{\prime}(t)\right)+\textsf{tr}\left(\nabla f(\gamma(t))^{\top}\gamma^{\prime\prime}(t)\right).\end{split}$ (38)
@@ A.2 Proofs of the results from Section 3
Proof of Lemma 3.1.
Let $\phi_{R}:A\in(\real^{p_{1}\times p_{2}})_{*}^{p}\rightarrow[A_{1},\ldots,A_{p}]\in\real^{p_{1}\times pp_{2}}$ and $\phi_{C}:A\in(\real^{p_{1}\times p_{2}})_{*}^{p}\rightarrow[A_{1}^{\top},\ldots,A_{p}^{\top}]\in\real^{p_{2}\times pp_{1}}$. For any $A\in(\real^{p_{1}\times p_{2}})_{*}^{p}$, $A_{R}\equiv\phi_{R}(A)\phi_{R}(A)^{\top}$ and $A_{C}\equiv\phi_{C}(A)\phi_{C}(A)^{\top}$ are of full-rank if and only if $\phi_{R}(A)$ and $\phi_{C}(A)$ are so, respectively. Thus, $\mathcal{H}_{p_{1},p_{2}}=\phi_{R}^{-1}(\real_{*}^{p_{1}\times pp_{2}})\cap\phi_{C}^{-1}(\real_{*}^{p_{2}\times pp_{1}})$. Note that both the maps $\phi_{R}$ and $\phi_{C}$ are smooth, and $\real{}_{*}^{p_{1}\times pp_{2}}$ and $\real{}_{*}^{p_{2}\times pp_{1}}$ are open in their respective ambient space. Thus, both $\phi_{R}^{-1}(\real_{*}^{p_{1}\times pp_{2}})$ and $\phi_{C}^{-1}(\real_{*}^{p_{2}\times pp_{1}})$ are open in $(\real^{p_{1}\times p_{2}})_{*}^{p}$. Hence, the transversality theorem ([38], Theorem 6.35) implies that $\mathcal{H}_{p_{1},p_{2}}$ is an open submanifold of $(\real^{p_{1}\times p_{2}})_{*}^{p}$. Because $(\real^{p_{1}\times p_{2}})_{*}^{p}$ is diffeomorphic to $\real{}_{*}^{p\times p}$ via the map $\varphi_{p_{1},p_{2}}$, $(\real^{p_{1}\times p_{2}})_{*}^{p}$ is open in $(\real^{p_{1}\times p_{2}})^{p}$. Therefore, $T_{A}\mathcal{H}_{p_{1},p_{2}}=T_{A}(\real^{p_{1}\times p_{2}})_{*}^{p}\equiv(\real^{p_{1}\times p_{2}})^{p}$ for any $A\in\mathcal{H}_{p_{1},p_{2}}$. ∎
Proof of Proposition 3.2.
Following the main idea outlined in the sketch of proof, it suffices to verify (15). Take $B=(B_{1},\ldots,B_{p})\in T_{A}\mathcal{H}_{p_{1},p_{2}}$. Note that $a^{\top}a=p$, $A_{R}=p_{2}I_{p_{1}}$, and $A_{C}=p_{1}I_{p_{2}}$. By Theorem $3.1$ of [44], $K_{(q,q)}$ is a symmetric matrix whose eigenvalues are either $1$ or 1-1, with respective multiplicities $q(q+1)/2$ and q(q1)/2q(q-1)/2. Moreover, the eigenspace of $K_{(q,q)}$ corresponding to the eigenvalue $1$ (resp. 1-1) is exactly the vectorization of $\mathcal{S}_{q}$ (resp. $\textsf{Skew}_{q}$). By vec-Kronecker identity,
@@
$(V-\textsf{tr}\left(V\right)/p_{1}I_{p_{1}})A_{i}=A_{i}(W-\textsf{tr}\left(W\right)/p_{2}I_{p_{2}}).$ (39)
@@
$\tilde{\Lambda}_{V}\tilde{A}_{i}=\tilde{A}_{i}\tilde{\Lambda}_{W}$ (40)
where $\tilde{\Lambda}_{V}=\Lambda_{V}-\textsf{tr}\left(\Lambda_{V}\right)/p_{1}I_{p_{1}}$ and $\tilde{\Lambda}_{W}=\Lambda_{W}-\textsf{tr}\left(\Lambda_{W}\right)/p_{2}I_{p_{2}}$. The equation (40) holds for each $i$ if and only if $((\tilde{\Lambda}_{V})_{jj}-(\tilde{\Lambda}_{W})_{kk})(\tilde{A}_{i})_{jk}=0$ for any $(j,k)\in[p_{1}]\times[p_{2}]$ and $i\in[p]$. Note that $\tilde{A}_{1},\ldots,\tilde{A}_{p}$ are linearly independent by the definition of $\mathcal{H}_{p_{1},p_{2}}$. Thus, for any fixed $(j,k)\in[p_{1}]\times[p_{2}]$, there must exist at least one $i\in[p]$ for which $(\tilde{A}_{i})_{jk}$ is nonzero, otherwise the linear independence is violated. Hence, $(\tilde{\Lambda}_{V})_{jj}=(\tilde{\Lambda}_{W})_{kk}$ for any $j,k$. As such, the diagonal entries of $\tilde{\Lambda}_{V}$ and $\tilde{\Lambda}_{W}$ are all equal to some constant $c$. However, as the traces of these matrices are the same as 0, both $\tilde{\Lambda}_{V}$ and $\tilde{\Lambda}_{W}$ are zero diagonal matrices so that $V=\Lambda_{V}=c_{1}I_{p_{1}}$ and $W=\Lambda_{W}=c_{2}I_{p_{2}}$ for some constants $c_{1},c_{2}$, leading to zero vectors $J_{1}^{\top}v$ and $J_{2}^{\top}w$. Hence, $C(J_{1}^{\top})\cap C(J_{2}^{\top})=\left\{0\right\}$, concluding the proof. ∎
Proof of Lemma 3.3.
where the last equality holds as $(X_{1},\ldots,X_{p})\in\mathcal{D}_{p_{1},p_{2}}$. Similarly, $\tilde{Y}_{C}\tilde{Y}_{C}^{\top}=p_{1}I_{p_{2}}$. ∎
Proof of Lemma 3.4.
By the quotient manifold theorem ([38], Theorem $21.10$), both $\mathcal{M}/G$ and $\mathcal{N}/G$ are smooth manifolds. Let $i:\mathcal{N}/G\xhookrightarrow{}\mathcal{M}/G$ be an inclusion map. We claim that $i$ is a smooth embedding. By Theorem $4.4$ of [71], $i$ is an injective immersion. To show that $i$ is a topological embedding, let $\pi:X\in\mathcal{M}\rightarrow[X]\in\mathcal{M}/G$ be the canonical projection, which is a smooth submersion. Hence, $\pi$ is open and $\mathcal{N}$ is $G$-invariant so that $\mathcal{N}$ is saturated for the map $\pi$. Here the subset $C\subset U$ is saturated for the map $f:U\rightarrow V$ between two topological spaces if $f^{-1}(f(C))=C$. Since $\mathcal{N}$ is a topological subspace of $\mathcal{M}$ as it is embedded, so is $\pi(\mathcal{N})\equiv\mathcal{N}/G$ of the quotient space $\pi(\mathcal{M})\equiv\mathcal{M}/G$ ([50], Theorem $22.1$). Thus, $i$ is a topological embedding. ∎
Proof of Proposition 3.6.
$\textsf{tr}_{1}(W)=\mathbf{0}_{p_{1}\times p_{1}},\quad\textsf{tr}_{2}(W)=\mathbf{0}_{p_{2}\times p_{2}}.$ (41)
@@
as $r>\alpha/\beta+\beta/\alpha\geq 2$ and $\alpha,\beta>2$, we conclude that the maximum of $m$ is strictly smaller than $r\alpha\beta$. ∎
Proof of Proposition 4.3.
@@
$\mathcal{P}_{U_{1}}A_{i}(I-\mathcal{P}_{W_{1}})=0,\quad(I-\mathcal{P}_{U_{1}})A_{i}\mathcal{P}_{W_{1}}=0.$ (42)
Recall that the orthogonal projection into a $d$-dimensional linear subspace of n corresponds to a unique element in a projective variety $\text{Gr}(d,n)$ as every linear subspace has its unique orthogonal projection. Namely, suppose $Z$ denotes a $d\times n$ matrix whose rows denote the basis of a $d$-dimensional subspace. Then the orthogonal projection onto this subspace is given by $Z^{\top}\textsf{adj}(ZZ^{\top})Z/|ZZ^{\top}|$, where $\textsf{adj}(M)$ is an adjoint of a square matrix $M$. By Theorem $2.1$ of [20], each entry of $Z^{\top}\textsf{adj}(ZZ^{\top})Z$ and $|ZZ^{\top}|$ can be expressed as a quadratic polynomial in Plüker coordinates. Note that $|ZZ^{\top}|\neq 0$. Hence, multiplying the quadratic polynomial corresponding to $|ZZ^{\top}|$ in Plüker coordinates to both hand sides of the equations in (42), we see that (42) induces the system of finitely many equations of polynomials in Plüker coordinates and affine coordinates (the entries of $A_{1},\ldots,A_{r}$). Hence, if $\tilde{\mathcal{V}}_{p_{1},p_{2},r}^{(a,b)}$ is a subset of $\mathcal{Z}:=\real\mathbb{P}^{\binom{p_{1}}{a}-1}\times\real\mathbb{P}^{\binom{p_{2}}{b}-1}\times\real^{pr}$ for which (42) is satisfied, then $\tilde{\mathcal{V}}_{p_{1},p_{2},r}^{(a,b)}$ is Zariski-closed as it is exactly the zero set of finitely many polynomials over $\mathcal{Z}$. Furthermore, since there are $ab+(p_{1}-a)(p_{2}-b)$ free coordinates for each $A_{i}$ with fixed $\mathcal{P}_{U_{1}},\mathcal{P}_{W_{1}}$,